# Speaking at the Right Level: Literacy-Controlled Counterspeech Generation with RAG-RL

### v2

@@ Appendix C Knowledge Bases
We detail the knowledge bases designed for different health literacy levels in Table LABEL:tab:knowledge-bases. We collect knowledge from a diverse set of reliable sources to ensure inclusiveness and representativeness across user groups, referring to previous studies Song et al. (2025a); Hong et al. (2025); Anik et al. (2025). These sources include federal health agencies such as the Centers for Disease Control and Prevention (CDC), academic and research institutions like the Johns Hopkins Children’s Center and the Allen Institute for AI, as well as community organizations including the Maryland Developmental Disabilities Council, St. Clair County Community Mental Health (SCCCMH), and the Independent Living Center of the Hudson Valley. These organizations provide materials written in plain language, often tailored to individuals with diverse health literacy backgrounds.
@@ Appendix E Evaluation Validation
Considering that user preference evaluations rely on LLMs, which may introduce potential bias and may not fully capture real users’ perspectives, we additionally conduct a human evaluation. We recruit six annotators representing low, medium, and high health literacy levels. To select annotators with diverse health literacy, we administer a brief screening survey following the method proposed by Rasmussen et al. (2023). Using this method, respondents rate each item on a 4-point scale using the following categories: Not at all true (1 point), Not completely true (2 points), Somewhat true (3 points), and Absolutely true (4 points). A cumulative score is then calculated. Based on the scoring criteria from the HLSAC, the initial version of the B-HLA categorizes health literacy as follows: low health literacy (10–25 points), moderate health literacy (26–35 points), and high health literacy (36–40 points). Two annotators are selected from each health literacy category, resulting in a balanced sample across literacy levels. All annotators are proficient in English and have no prior involvement in the project. Annotators are compensated at a rate of \$10 per hour, in accordance with ethical guidelines for human-subject research. Each annotator completes the evaluation in approximately two hours. The final annotator group consisted of individuals aged 15 to 35 years, with educational backgrounds ranging from high school to graduate-level studies. The group includes three female and three male participants, all based in the United States. All annotators reported regular access to online health information, though their confidence and comprehension varied, as reflected in the screening tool scores.
@@
We present the results in Table LABEL:tab:correlation. The results show that Pearson correlation is highest for low and medium literacy levels (0.67–0.68), suggesting the LLM matches human scores most closely in absolute value for users with lower literacy. Spearman and Kendall’s Tau are highest at the high literacy level, indicating that the LLM is especially good at ranking outputs in the same order as human annotators at this level, even if the exact scores differ. These findings suggest that the LLM’s behavior adapts well to different user profiles: (1) For low and medium literacy users, the focus is on numeric simplicity and readability, and LLM-generated scores align well with human ratings in both value and order. (2) For high literacy users, while the Pearson correlation is slightly lower, the LLM captures the ranking preferences of more sophisticated users very effectively.
@@ Appendix F Top-$k$ Comparison
We use the LLaMA3.1-8B-Instruct model within a RAG framework to examine the impact of Top-$k$ evidence selection. Results in Table LABEL:tab:topk_eval reveal the following: for users with low health literacy, Top-5 achieves the highest politeness score (0.73) and lowest target distance (1.14), Top-3 yields the highest user preference score (0.73), and Top-10 leads in factual accuracy (0.89). Given the relatively small differences in politeness, user preference, and target distance but a more substantial advantage in factual accuracy, we select Top-10 for users with low health literacy. For users with medium health literacy, Top-3 achieves the best overall performance across all dimensions: politeness (0.71), target distance (4.45), user preference (0.73), and factual accuracy (0.88), and is therefore selected. In the high health literacy setting, Top-10 is chosen as it demonstrates the best performance in user preference (0.72) and factual accuracy (0.90).
@@ Appendix G Cross Generalization Results
We apply our methods, Controlled-Literacy, to two distinct misinformation datasets: Check-COVID and MisinfoCorrect mentioned in Section 5. The results are in Table LABEL:tab:checkcovid_results and Table LABEL:tab:misinfo_correct_results.

### v3

@@ Appendix C Knowledge Bases
We detail the knowledge bases designed for different health literacy levels in Table 6. We collect knowledge from a diverse set of reliable sources to ensure inclusiveness and representativeness across user groups, referring to previous studies Song et al. (2025a); Hong et al. (2025); Anik et al. (2025). These sources include federal health agencies such as the Centers for Disease Control and Prevention (CDC), academic and research institutions like the Johns Hopkins Children’s Center and the Allen Institute for AI, as well as community organizations including the Maryland Developmental Disabilities Council, St. Clair County Community Mental Health (SCCCMH), and the Independent Living Center of the Hudson Valley. These organizations provide materials written in plain language, often tailored to individuals with diverse health literacy backgrounds.
@@ D.1 Instructional Prompt
Health Literacy Level Top-$k$ Politeness Target Distance User Preference Factual Accuracy
Low top_10 0.72 (0.25) 1.30 (3.23) 0.71 (0.13) 0.89
top_5 0.73 (0.25) 1.14 (2.97) 0.71 (0.12) 0.87
top_3 0.71 (0.25) 1.23 (3.14) 0.73 (0.09) 0.84
Medium top_10 0.66 (0.22) 4.93 (6.81) 0.71 (0.13) 0.81
top_5 0.70 (0.22) 4.98 (6.76) 0.72 (0.10) 0.87
top_3 0.71 (0.23) 4.45 (6.35) 0.73 (0.11) 0.88
High top_10 0.41 (0.21) 0.08 (0.75) 0.72 (0.09) 0.90
top_5 0.43 (0.21) 0.04 (0.72) 0.71 (0.10) 0.89
top_3 0.42 (0.21) 0.42 (2.56) 0.71 (0.10) 0.81
Table 7: Comparison of Top-$k$ selections.
@@ D.3 Evaluation Prompt - Factual Accuracy
Method Literacy Level Politeness Target Distance (↓) User Preference Factual Accuracy
LLaMA-8B
Instructional Prompt low 0.52 (0.19) 2.69 (4.25) 0.74 (0.06) 0.87
medium 0.60 (0.17) 4.94 (7.49) 0.74 (0.08) 0.86
high 0.51 (0.12) 0.03 (0.56) 0.74 (0.07) 0.76
Avg. 0.54 (0.16) 2.55(4.10) 0.74 (0.07) 0.83
RAG low 0.72 (0.19) 2.11 (4.06) 0.74 (0.06) 0.82
medium 0.74 (0.16) 3.21 (5.82) 0.74 (0.07) 0.85
high 0.57 (0.13) 0.05 (0.60) 0.70 (0.11) 0.88
Avg. 0.68 (0.16) 1.79 (3.49) 0.73 (0.08) 0.85
Controlled-Literacy low 0.82 (0.12) 0.74 (1.90) 0.75 (0.00) 0.92
medium 0.75 (0.15) 2.40 (4.26) 0.75 (0.02) 0.91
high 0.99 (0.00) 0.00 (0.00) 0.75 (0.02) 0.96
Avg. 0.85 (0.09) 1.05 (2.05) 0.75 (0.01) 0.93
LLaMA-1B
Instructional Prompt low 0.68 (0.21) 4.36 (7.40) 0.65 (0.19) 0.57
medium 0.50 (0.19) 17.29 (11.74) 0.55 (0.23) 0.59
high 0.84 (0.17) 0.06 (0.60) 0.65 (0.16) 0.46
Avg. 0.67 (0.19) 7.24 (6.58) 0.62 (0.19) 0.54
RAG low 0.52 (0.19) 16.90 (14.68) 0.41 (0.24) 0.69
medium 0.55 (0.15) 9.19 (10.50) 0.43 (0.23) 0.59
high 0.57 (0.17) 0.11 (1.27) 0.43 (0.21) 0.51
Avg. 0.55 (0.17) 8.73 (8.82) 0.42 (0.23) 0.60
Controlled-Literacy low 0.80 (0.14) 2.29 (3.69) 0.71 (0.13) 0.69
medium 0.79 (0.17) 2.22 (3.81) 0.72 (0.10) 0.74
high 0.83 (0.27) 0.05 (0.73) 0.68 (0.13) 0.75
Avg. 0.81 (0.19) 1.52 (2.74) 0.70 (0.12) 0.73
Qwen-7B
Instructional Prompt low 0.46 (0.18) 4.93 (7.12) 0.73 (0.10) 0.84
medium 0.59 (0.18) 4.79 (7.69) 0.74 (0.11) 0.95
high 0.53 (0.13) 0.58 (6.05) 0.74 (0.10) 0.97
Avg. 0.53 (0.16) 3.43 (6.95) 0.74 (0.10) 0.92
RAG low 0.58 (0.17) 5.20 (6.98) 0.74 (0.07) 0.88
medium 0.64 (0.16) 7.09 (8.22) 0.75 (0.04) 0.95
high 0.63 (0.14) 0.03 (0.38) 0.73 (0.08) 0.92
Avg. 0.62 (0.16) 4.11 (5.19) 0.74 (0.06) 0.92
Controlled-Literacy low 0.79 (0.16) 1.94 (3.71) 0.75 (0.04) 0.90
medium 0.62 (0.13) 1.36 (2.67) 0.75 (0.04) 0.97
high 0.88 (0.18) 0.00 (0.00) 0.74 (0.05) 0.98
Avg. 0.76 (0.16) 1.10 (2.13) 0.75 (0.04) 0.95
Table 8: Cross generalization performance on CheckCovid Dataset. The best overall performance in each category is highlighted in gray.
@@ Appendix E Evaluation Validation
Considering that user preference evaluations rely on LLMs, which may introduce potential bias and may not fully capture real users’ perspectives, we additionally conduct a human evaluation. We recruit six annotators representing low, medium, and high health literacy levels. To select annotators with diverse health literacy, we administer a brief screening survey following the method proposed by Rasmussen et al. (2023). Using this method, respondents rate each item on a 4-point scale using the following categories: Not at all true (1 point), Not completely true (2 points), Somewhat true (3 points), and Absolutely true (4 points). A cumulative score is then calculated. Based on the scoring criteria from the HLSAC, the initial version of the B-HLA categorizes health literacy as follows: low health literacy (10–25 points), moderate health literacy (26–35 points), and high health literacy (36–40 points).
Method Literacy Level Politeness Target Distance (↓) User Preference Factual Accuracy
LLaMA-8B
Instructional Prompt low 0.41 (0.24) 2.17 (4.03) 0.75 (0.02) 0.98
medium 0.42 (0.20) 2.94 (5.59) 0.75 (0.01) 0.98
high 0.29 (0.17) 0.11 (0.93) 0.75 (0.04) 0.99
Avg. 0.37 (0.20) 1.74 (3.52) 0.75 (0.02) 0.98
RAG low 0.86 (0.15) 0.98 (2.55) 0.75 (0.00) 0.99
medium 0.72 (0.24) 1.49 (2.84) 0.75 (0.03) 1.00
high 0.34 (0.18) 0.13 (1.07) 0.74 (0.04) 0.97
Avg. 0.64 (0.19) 0.87 (2.15) 0.75 (0.02) 0.99
Controlled-Literacy low 0.92 (0.09) 0.78 (1.95) 0.75 (0.00) 1.00
medium 0.68 (0.23) 0.89 (2.45) 0.75 (0.00) 1.00
high 0.97 (0.02) 0.00 (0.00) 0.75 (0.00) 1.00
Avg. 0.86 (0.11) 0.56 (1.47) 0.75 (0.00) 1.00
LLaMA-1B
Instructional Prompt low 0.73 (0.28) 12.41 (12.03) 0.64 (0.18) 0.62
medium 0.47 (0.32) 15.82 (16.05) 0.59 (0.22) 0.72
high 0.82 (0.18) 0.14 (2.54) 0.68 (0.15) 0.78
Avg. 0.67 (0.26) 9.46 (10.21) 0.64 (0.18) 0.71
RAG low 0.43 (0.30) 15.53 (12.60) 0.53 (0.23) 0.68
medium 0.36 (0.25) 5.59 (7.89) 0.48 (0.23) 0.79
high 0.33 (0.23) 0.51 (3.78) 0.49 (0.21) 0.76
Avg. 0.37 (0.26) 7.21 (8.09) 0.41 (0.22) 0.74
Controlled-Literacy low 0.84 (0.16) 4.54 (5.18) 0.68 (0.17) 0.74
medium 0.61 (0.23) 1.05 (2.30) 0.72 (0.08) 0.70
high 0.92 (0.20) 0.00 (0.00) 0.68 (0.13) 0.81
Avg. 0.79 (0.20) 1.86 (2.49) 0.69 (0.13) 0.75
Qwen-7B
Instructional Prompt low 0.31 (0.22) 3.04 (5.74) 0.74 (0.06) 0.91
medium 0.45 (0.23) 4.74 (10.06) 0.74 (0.10) 0.97
high 0.38 (0.21) 0.02 (0.41) 0.74 (0.05) 1.00
Avg. 0.38 (0.22) 2.60 (5.40) 0.74 (0.07) 0.96
RAG low 0.59 (0.22) 3.96 (6.38) 0.75 (0.01) 0.97
medium 0.52 (0.22) 4.84 (7.15) 0.75 (0.03) 0.96
high 0.49 (0.26) 0.02 (0.17) 0.74 (0.04) 1.00
Avg. 0.53 (0.23) 2.94 (4.57) 0.75 (0.03) 0.98
Controlled-Literacy low 0.78 (0.19) 2.32 (2.99) 0.75 (0.00) 1.00
medium 0.43 (0.16) 1.23 (2.53) 0.75 (0.03) 0.99
high 0.77 (0.30) 0.00 (0.00) 0.75 (0.00) 1.00
Avg. 0.66 (0.22) 1.18 (1.84) 0.75 (0.01) 1.00
Table 9: Cross generalization performance on MisinfoCorrect Dataset. The best overall performance in each category is highlighted in gray.
Two annotators are selected from each health literacy category, resulting in a balanced sample across literacy levels. All annotators are proficient in English and have no prior involvement in the project. Annotators are compensated at a rate of \$10 per hour, in accordance with ethical guidelines for human-subject research. Each annotator completes the evaluation in approximately two hours. The final annotator group consisted of individuals aged 15 to 35 years, with educational backgrounds ranging from high school to graduate-level studies. The group includes three female and three male participants, all based in the United States. All annotators reported regular access to online health information, though their confidence and comprehension varied, as reflected in the screening tool scores.
@@
We present the results in Table 3. The results show that Pearson correlation is highest for low and medium literacy levels (0.67–0.68), suggesting the LLM matches human scores most closely in absolute value for users with lower literacy. Spearman and Kendall’s Tau are highest at the high literacy level, indicating that the LLM is especially good at ranking outputs in the same order as human annotators at this level, even if the exact scores differ. These findings suggest that the LLM’s behavior adapts well to different user profiles: (1) For low and medium literacy users, the focus is on numeric simplicity and readability, and LLM-generated scores align well with human ratings in both value and order. (2) For high literacy users, while the Pearson correlation is slightly lower, the LLM captures the ranking preferences of more sophisticated users very effectively.
@@ Appendix F Top-$k$ Comparison
We use the LLaMA3.1-8B-Instruct model within a RAG framework to examine the impact of Top-$k$ evidence selection. Results in Table 7 reveal the following: for users with low health literacy, Top-5 achieves the highest politeness score (0.73) and lowest target distance (1.14), Top-3 yields the highest user preference score (0.73), and Top-10 leads in factual accuracy (0.89). Given the relatively small differences in politeness, user preference, and target distance but a more substantial advantage in factual accuracy, we select Top-10 for users with low health literacy. For users with medium health literacy, Top-3 achieves the best overall performance across all dimensions: politeness (0.71), target distance (4.45), user preference (0.73), and factual accuracy (0.88), and is therefore selected. In the high health literacy setting, Top-10 is chosen as it demonstrates the best performance in user preference (0.72) and factual accuracy (0.90).
@@ Appendix G Cross Generalization Results
We apply our methods, Controlled-Literacy, to two distinct misinformation datasets: Check-COVID and MisinfoCorrect mentioned in Section 5. The results are in Table 8 and Table 9.